Steven E Jens

20 Lee Street Cambridge, MA 02139 steven@stevenjens.com 617-864-3311

Objective

An entry-level position in investment management.

Education

BU Graduate School of Management Boston, Massachusetts

Master’s of Science in Investment Management cum laude (2004).

Cumulative GPA: 3.61.

GMAT: 780 The Massachusetts Institute of Technology Cambridge, Massachusetts

BS in Mathematics with Computer Science with minors in Economics and Mechanical Engineering (1998).

Experience

Foresight Research Solutions, Intern 2004

Updated and maintained Excel-based earnings models for exploration and production companies in the energy industry.

USPowerSolutions LLC, Java Developer 1998-2001

Wrote software to assist energy suppliers in communication with utility companies. Used Java, including EJBs (on WebLogic) and JSP.

InterSystems Corporation, Object group 1997-1998 (part-time during the school year)

Wrote and tested a feature to convert M/MUMPS (a programming language produced by InterSystems) objects into Java.

Athena Consulting, MIT Information Services 1996-1997

Resolved user questions about MIT’s main computer network and the software thereon. Advised students about advanced computer systems and software.

Skills

Financial Skills/Background

Passed the CFA Level II Exam. Computer Skills

Quite familiar with Excel. Unix (Solaris and Linux), HTML, emacs, basic makefiles, X12 EDI, and various Internet protocols.

Languages include Java, Perl, Postscript, C, Excel-VBA, C++, and Emacs Lisp. Languages

Speak intermediate French. City Council candidate (2001)

Organized a team of about ten volunteers, seventy contributors, and several hundred voters, placing sixth of twelve challengers.

http://www.stevenjens.com/resume/